OpenStudy (anonymous):

What is 20% of $29.99

OpenStudy (anonymous):

20*29.99/100 = 5.998

OpenStudy (shamil98):

\[\huge \frac{ part }{ whole } = \frac{ \text{%} }{ 100 }\] Use that.

OpenStudy (anonymous):

5.998